OpenStudy (anonymous):

what is 24 1/4 take away 5 5/6

OpenStudy (anonymous):

\[24\frac{ 1 }{ 4 }-5\frac{ 5 }{ 6 } = ?\]

OpenStudy (anonymous):

Take both the mixed fractions as improper fractions, then subract as you normally would.

OpenStudy (anonymous):

how do u answer it

OpenStudy (anonymous):

To convert 24 1/4 into an improper fraction, multiply 24 by 4 then add 1 to it and that's your numerator divided by 4 which is the denominator. Same for 5 5/6, then subtract them normally. \[\frac{ 97 }{ 4 }-\frac{ 35 }{ 6 } = ?\]

OpenStudy (anonymous):

Take L.C.M. of the denominators, and then subtract.

OpenStudy (anonymous):

huh